OpenStudy (anonymous):

The volume of a cylinder is 93 ft3. The volume of a cube is 48 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube? Provide an explanation for your answer to receive full credit. Can anyone please help me with this question?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@bahrom7893 can u plz help?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@agent0smith

jimthompson5910 (jim_thompson5910):

Hints: Volume of Cone = (1/3)*(Volume of Cylinder) Note: both the cone and cylinder have the same radius and height -------------------- Volume of Pyramid = (1/3)*(Volume of Cube) Note: both the pyramid and cube have the same dimensions

OpenStudy (anonymous):

so what do i do now?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

"The volume of a cylinder is 93 ft^3" so the volume of the cone is ???

OpenStudy (anonymous):

idk..

OpenStudy (anonymous):

help? hints?

jimthompson5910 (jim_thompson5910):

use Volume of Cone = (1/3)*(Volume of Cylinder)

OpenStudy (anonymous):

31 ft^3?..

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

"The volume of a cube is 48 ft^3" volume of pyramid = ???

OpenStudy (anonymous):

16 ft^3?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

now find the difference of the cone and pyramid

OpenStudy (anonymous):

volume of a pyramid is 16 ft^3 and volume of a cone is 31 ft^3

jimthompson5910 (jim_thompson5910):

subtract the two now

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so.. how do we get to our answer?

jimthompson5910 (jim_thompson5910):

you just got it

OpenStudy (anonymous):

so how do i put all of this together?

jimthompson5910 (jim_thompson5910):

show how to get the volumes of the cone and pyramid then show the subtraction to show the difference in the two volumes

OpenStudy (anonymous):

so i say to get the volume of the cone i had to multiply 93 by 1/3 and to get the volume of the pyramid i had to multiply 48 by 1/3. then i just subtracted 16 from 31 to get 15 as my answer?..

OpenStudy (anonymous):

@jim_thompson5910